OpenStudy (anonymous):

why does 2^n(n+1)+2^n(n-1) simplify to n2^(n+1)??? Please help :)

myininaya (myininaya):

Is that \[2^{n(n+1)}+2^{n(n-1)} ?\]

OpenStudy (anonymous):

sorry, \[2^{n}(n+1)+2^{n}(n-1)\]

myininaya (myininaya):

\[2^n(n+1+n-1)= ?\]

OpenStudy (anonymous):

the anwer would be\[ n2^{n+1}\] but I don't understand how... :P

myininaya (myininaya):

what is n+1+n-1?

myininaya (myininaya):

\[2^n(2n)=2^n \cdot 2 \cdot n =2^n \cdot 2^1 \cdot n\] By law of exponents we can do something with the 2^n times 2^1

OpenStudy (anonymous):

Ohhhh! I get it. the ones cancel out in 2^n(n+1+n−1) becoming 2n. Thank you very much!!!

OpenStudy (anonymous):

woah, so many ?

myininaya (myininaya):

reload

myininaya (myininaya):

they go away after reloading lol

OpenStudy (anonymous):

Ah thank you very much!

myininaya (myininaya):

np